OpenStudy (anonymous):

(2, -3) is a solution of the equation 3x - y = 3.

OpenStudy (anonymous):

Hey This is almost the same question of the previous one u answered. please try to work it out yourself. In all honesty, it is not a difficult one.

OpenStudy (anonymous):

asked* (not answered)

OpenStudy (anonymous):

Dont know why u erased your answer, You got it right. 3x-y=3 and (2,-3) Substitute, 3(2) - (-3) = 3 6+3 = 3 9=3 Which is wrong. Thus not a solution. Be confident of your answers.

OpenStudy (anonymous):

Is (-1, -1) a solution of y = 3x + 2 -1=-1 so its true

OpenStudy (anonymous):

Yes! You got that right!